OpenStudy (anonymous):

I need help with x+2 divided by x-8

OpenStudy (unklerhaukus):

change the numerator a bit\[\frac{x+2}{x-8}=\frac{x-8+2+8}{x-8}=\frac{x-8}{x-8}+\frac{2+8}{x-8}\] can you see what i have done?

OpenStudy (anonymous):

no

OpenStudy (unklerhaukus):

i have taken away 8 from the numerator to get something that will cancel, but to keep the fraction equal i needed to add 8 as well, now the fraction has two parts that can be evaluated separately, the first one cancels nicely

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answer is 2 over x

OpenStudy (unklerhaukus):

how did you get that?

OpenStudy (anonymous):

i cross out the 8's

OpenStudy (anonymous):

sorry math is my worst subject i'm trying to learn this algerba

OpenStudy (unklerhaukus):

\[\frac{x-8}{x-8}+\frac{2+8}{x-8}\] the first fraction will cancel , but for the second one all you can do is to simplify the numerator

OpenStudy (unklerhaukus):

what do you get?

OpenStudy (unklerhaukus):

What exactly was the original question?

OpenStudy (unklerhaukus):

oh , now i understand the question, i thought it was something different, lets start again.

OpenStudy (unklerhaukus):

\[\frac{x+2}{x-8}\] What you need to know is that a fraction wont be defined when the denominator is zero so what value(s) of x make x-8 equal to zero?

OpenStudy (anonymous):

8-8=0

OpenStudy (unklerhaukus):

yeah when x is 8, the denominator becomes zero, and the fraction wont make sense, any number divided by zero is undefined

OpenStudy (unklerhaukus):

well i would just set denominator to 0 and solve for x x-8=0 x-8+8=8 x=8

OpenStudy (anonymous):

thank you one more problem write in lowest term a2-b2 divided by b-a

OpenStudy (anonymous):

that is a square minus b square

OpenStudy (unklerhaukus):

\[\frac{a^2-b^2}{b-a}\]

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

write the lowest term? is that really the whole question?

OpenStudy (anonymous):

yes the answer I got is negative 1. just want to make sure it is right.

OpenStudy (unklerhaukus):

is there some more information? i still dont understand this question

OpenStudy (anonymous):

a squared minus b squared divided by a minus b

OpenStudy (unklerhaukus):

how did you get -1?

OpenStudy (anonymous):

I crossed out the each a's and b's and when I cross them out i gave each a one and diveded negative l by one.

OpenStudy (anonymous):

i'm confused, the problem is( a squared minus b squared )divided by (a minus b) yes the bootom does not squared so Im not able to cross out is that what you are saying.

OpenStudy (unklerhaukus):

you can only cancel terms in fractions when they are factors, (ie with multiplication, [not addition or subtraction])

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

Cam you tell me the exact wording of this question , i think im still missing part of it

OpenStudy (anonymous):

it only says to write in lowest terms.

OpenStudy (unklerhaukus):

Oh, "Write the expression \(\frac{a^2-b^2}{b-a}\) in lowest terms"

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

This just means you have to expand the numerator using difference of squares a^2−b^2=(a−b)(a+b) and then cancel the common factor (after taking the negative, to get the factor in the numerator the same as the denominator

OpenStudy (anonymous):

how do I do that?

OpenStudy (unklerhaukus):

first re-write the numerator as a product of two factors (as i have above,)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you so much for youe help. If I need any more help with equations I hope that I can adrress them to you if noone else response once again thanks.

OpenStudy (unklerhaukus):

yeah for next time , i will try and help for sure, but now i have to go to sleep bye

OpenStudy (anonymous):

bye